$SPAD/src/input explot2d.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input for page ExPlot2DPolar
)clear all
draw(sin(4*t/7),t = 0..14*%pi,coordinates == polar)
-- Input for page ExPlot2DAlgebraic
)clear all
draw(solution(y**2 + y - (x**3 - x) = 0),x,y,clip == (-2..2,-2..1))
-- Input for page ExPlot2DFunctions
)clear all
draw(sin(tan(x)) - tan(sin(x)),x = 0..6)
-- Input for page ExPlot2DParametric
)clear all
draw(curve(9 * sin(3*t/4),8 * sin(t)),t = -4*%pi..4*%pi)

———-

References^{[1] nothing}

4